OpenStudy (anonymous):

How do you use the reverse power rule? f(x)=5x^4-4x^3+3x^2-2x+1

OpenStudy (anonymous):

add one to the exponent and divide by that number \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

how do i apply that to this problem?

OpenStudy (anonymous):

piece by piece

OpenStudy (anonymous):

can you give an example? *right now i'm just studying, so use can use the problem above i made up*

OpenStudy (anonymous):

\[\int 5x^4dx=5\times\frac{x^{4+1}}{4+1}\]

OpenStudy (anonymous):

i.e. \[5\frac{x^5}{5}=x^5\]

OpenStudy (freckles):

\[f(x)=ax^2+bx+c \\ F(x)=a \frac{x^{2+1}}{2+1}+b \frac{x^{1+1}}{1+1}+cx+d\] Where F'=f

OpenStudy (freckles):

that was another example

OpenStudy (anonymous):

could you do 5x^4 step-by-step?

OpenStudy (anonymous):

i don't get this :(

OpenStudy (freckles):

\[\text{ if } g(x)=5x^4 \text{ then } G(x)=5 \cdot \frac{x^{4+1}}{4+1}+c\]

OpenStudy (freckles):

That is the absolute first step

OpenStudy (freckles):

\[f(x)=ax^{n} \text{ then } F(x)=a \cdot \frac{x^{n+1}}{n+1} +c \text{ where } n \neq -1 \]

OpenStudy (freckles):